### v1

# Cluster-Dags as Powerful Background Knowledge For Causal Discovery

Abstract
Finding cause-effect relationships is of key importance in science. Causal discovery aims to recover a graph from data that succinctly describes these cause-effect relationships. However, current methods face several challenges, especially when dealing with high-dimensional data and complex dependencies. Incorporating prior knowledge about the system can aid causal discovery. In this work, we leverage Cluster-DAGs as a prior knowledge framework to warm-start causal discovery. We show that Cluster-DAGs offer greater flexibility than existing approaches based on tiered background knowledge and introduce two modified constraint-based algorithms, Cluster-PC and Cluster-FCI, for causal discovery in the fully and partially observed setting, respectively. Empirical evaluation on simulated data demonstrates that Cluster-PC and Cluster-FCI outperform their respective baselines without prior knowledge.
1 Introduction
Understanding causal relationships is essential for scientific inquiry and reasoning about the world. Researchers have always leveraged active experimentation and interventions to uncover causal mechanisms. But in many cases, such experiments are impractical or ethically off-limits—for instance, it would be unethical to expose communities to varying levels of environmental pollution to study their effect on health. The challenges of experimentation, along with the data explosion in recent decades, have led to continued methodological advances in causal discovery from observational data (guo2020survey; malinsky2018causal; peters2017elements).
One common approach to represent causal assumptions is the Structural Causal Model (SCM) framework (pearl2009causal; peters2017elements), which encodes these relationships in Directed Acyclic Graphs (DAGs), with edges representing the “directed flow of causal influence” between variables. One landmark result in causal discovery is that, without further assumptions, purely from observational data one can only identify the Markov equivalence class—a collection of graphs that are all consistent with the observed data. In large graphs, this ambiguity can make it difficult to pinpoint precise causal insights (spirtes2000causation; spirtes1995causal). Moreover, as the number of variables increases, the space of DAGs grows exponentially, posing serious scalability issues for many causal discovery algorithms (ijcai2024p374).
One way to help address these challenges is by incorporating background knowledge—partial information about the graph structure that can narrow the search space and guide causal inference. A recent method for encoding such partial knowledge is via Cluster-DAGs (C-DAG) (anand2023causal), which organize variables into clusters and assume that possible causal relationships between these clusters are known, but the causal structure within the clusters as well as the precise connections of individual variables across clusters is still unknown. By using structural information at this higher level, C-DAGs offer a way to manage complexity in high-dimensional settings. C-DAGs were originally developed for direct use in causal inference, reasoning about the strength of cause-effect relationships using the C-DAG directly (anand2023causal).
We leverage C-DAGs as background knowledge to improve constraint-based causal discovery methods using the prior knowledge during discovery process itself. We also extensively compare it to the related tiered background knowledge proposed by (andrews2020completeness). Our contributions include:
  • We show that tiered background knowledge (TBK) (andrews2020completeness) can be represented as a C-DAG, but not vice-versa—C-DAGs capture strictly more flexible types of background knowledge. In addition, C-DAGs can accommodate latent confounding between clusters, whereas TBK has to assume it non-existent, see Section˜2.1.
  • We formalize the constraints a C-DAG puts on a corresponding DAG by formulating the C-DAG restrictions as a boolean combination of pairwise constraints, see Appendix˜C for details. This could be interesting for future theoretic research in causal discovery with background knowledge.
  • C-DAGs allow warm-starting of constraint-based causal discovery algorithms, namely PC (spirtes2000causation) and FCI (zhang2008completeness) by pruning and orienting edges before running the discovery algorithm. This results in fewer CI tests needing to be performed. We introduce the Cluster-PC and Cluster-FCI formally and show their soundness (C-PC and C-FCI) and completeness (for C-PC only, C-FCI is not be complete by design).
  • Our simulations show that Cluster-PC and Cluster-FCI outperform the non-background knowledge baselines, see Sections˜3 and 4. Cluster-FCI also outperforms FCITiers on C-ADMGs (allowing bidirected edges between clusters), while being close to FCITiers on C-DAGs.
1.1 Related work
There are two main ways in which background knowledge can be described in graphical structures: groupwise background knowledge and pairwise background knowledge. Pairwise background knowledge (fang2022representationcausalbackgroundknowledge; meek2013causalinferencecausalexplanation) restricts the relationship between pairs of variables, i.e., requiring or ruling out directed edges or ancestral relationships. In contrast, groupwise background knowledge (andrews2020completeness; brouillard2022typing; anand2023causal) organizes variables into different groups and then restricts the edges between these groups, without committing to any pairwise constraint directly.
Previous work on background knowledge includes guiding score-based methods like KGS (hasan2024optimizingdatadrivencausaldiscovery), which can use prior knowledge on the absence/presence of an (un)directed edge, $A^{*}$-based methods (kleinegesse2022domainknowledgeabasedcausal), using absence/ presence of directed edges and tiers (non-ancestral constraints) and NOTEARS (chowdhury2023evaluationinducedexpertknowledge), using absence/ presence of directed edges). Most of these methods focus on pairwise background knowledge. For constraint-based methods, except for (andrews2020completeness), background knowledge is usually used to orient additional edges after receiving the CPDAG (brouillard2022typing; bang2023wisertimecausalequivalence) from the PC algorithm (spirtes2000causation). fang2022representationcausalbackgroundknowledge study the representation of causal background knowledge using pairwise background knowledge. Pairwise and groupwise background knowledge can be combined, although to the best of our knowledge, not much research has gone in this direction yet.
Recently, different aspects of causal diagrams over groups of variables have been discussed. parviainen2017learning study learning groups of variables in Bayesian networks. wahl2023vector discuss two methods for inferring causal relationships between two groups of variables. melnychuk2024boundsrepresentationinducedconfoundingbias use C-DAGs to group confounders as a cluster. wahl2024foundations extend the theoretical framework of anand2023causal, discussing the relationships between micro and group level graphs. C-DAGs and other variable aggregation methods have increasingly been used to model causal systems (ma2025foundationmodelscausalinference; plecko2024mind; raghavan2025counterfactual; assaad2025identifiabilitytotaleffectsabstractions; xia2025causal; tabell2025clustering). Such an aggregate model can then, with our approach, be used to inform causal discovery of the more granular DAG.
li2024disentangled present a method for learning disentangled causal representations from very high-dimensional data like images. Group graph variants have also increasingly become targets for causal discovery (niulearning; ninad2025causaldiscoveryvectorvaluedvariables; gobler2025nonlinearcausaldiscoverygrouped; anand2025causal). xia2024neural use neural causal models to learn neural causal abstractions by clustering variables and their domains, while massidda2024learningcausalabstractionslinear create a variant of LiNGAM (shimizu2014lingam), Abs-LiNGAM, to learn causal abstractions. li2025identifiabilitycausalabstractions; yvernes2025identifiabilitycausalabstractionshierarchy investigate the identifiability of causal abstractions, schooltink2025aligninggraphicalfunctionalcausal bridge graphical and functional causal abstractions and massidda2025weakly study causal sufficiency for causal abstractions. While much work focuses on learning the aggregated group graph itself, our approach uniquely leverages the C-DAG as assumed prior knowledge to perform causal discovery that resolves the relationships within the underlying micro-variables.
The CaGreS algorithm by zeng2025causaldagsummarizationfull summarizes DAGs via node contractions and creates summary causal graphs (SCGs) with preserved utility for causal inference. ferreira2025identifyingC extend on anand2023causal by analyzing theoretical properties of SCGs, which are similar to C-DAGs, but also allow for cycles. yvernes2025completecharacterizationadjustmentsummary; assaad2025identifiabilitytotaleffectsabstractions; assaad2025identifiabilitymicrototaleffects; yvernes2025note; yvernes2025identifiabilitycommonbackdoorsummary; ferreira2025averagecontrolledaveragenatural investigate identifiability in SCGs, while ferreira2025identifyingmacrocausaleffects do so for cluster directed mixed graphs, a related concept. Transit clusters (tikka2023clustering), while similar, are specifically designed to cluster variables while preserving identifiability properties. zhu2024meaningful show that interventions in aggregation of variables (e.g., a surjective but non-injective function of cluster variables) are no longer well-defined. We build on these properties developed for C-DAGs to guide a constraint-based search for the detailed underlying DAG.
On the application side, C-DAGs have also shown some promise for improving causal inference tasks. ribeiro2025bites apply causal inference to assess malaria risk, and they mention that causal discovery at the cluster level improves interpretability. anand2025leveraging apply C-DAGs to determine causal effects in medicine.
Research gap: Despite increasing interest in causally modeling groups of variables, exploiting often easily available groupwise background knowledge for causal discovery of the detailed graph remains underexplored. Groupwise background knowledge flexibly accommodates prior knowledge on varying levels of detail for different subsets of the variables, often rendering it much more realistic in practice. Some groups of non-ancestrality constraints (e.g., today can not causally influence yesterday) is much more readily available and justified than individual required/forbidden edges or highly granular assertions. C-DAGs encode such groupwise knowledge flexibly and in a visually interpretable, intuitive way, making them a valuable tool for applied researchers and users across domains.
1.2 Preliminaries on causality and constraint-based causal discovery
We briefly introduce the most relevant preliminaries for structural causal models (SCM) and causal discovery. For more details, we refer the reader to (spirtes2000causation; pearl2009causality; peters2017elements). Throughout, we assume a fixed set of random variables $X_{1},\ldots,X_{n}$, which also serve as the node set of graphs $V=\{X_{1},\ldots,X_{n}\}$.
A DAG (directed acyclic graph) $G=(V,E)$ is a directed graph over nodes $V$ with edges $E$ without directed cycles. If $(X\rightarrow Y)\in E$ in a DAG (for $X,Y\in V$), we write $X\in pa_{Y}$, $Y\in ch_{X}$ (parents, children, respectively). The neighbors of $X$ are: $nb_{X}:=ch_{X}\cup pa_{X}$. A superscript $G$ like $an_{X}^{G}$ indicates we are referring to the ancestors of $X$ in graph $G$. If there exists a directed path from $X$ to $Y$, i.e., $X\rightarrow\ldots\rightarrow Y$, then $X\in an_{Y}$, $Y\in de_{X}$ (ancestors, descendants respectively). When we find $X\rightarrow Y\leftarrow Z$ along a path in $G$, we call $Y$ a collider on the path; if $X,Z$ are additionally not adjacent in the DAG, we call the triple $\{X,Y,Z\}\subset V$ an unshielded collider or v-structure. For causal discovery, we will also consider acyclic graphs that can contain both directed and undirected edges (denoted by XYX-Y). This will often be interpreted as the direction of the edge not yet being specified or unknown.
A structural causal model over variables in $V$ entails both a probability distribution $P(X_{1},\ldots,X_{n})$, the observational distribution, and a DAG $G=(V,E)$. In fully observed SCMs, the observational distribution and the implied DAG are related by the Markov property: conditional independence statements about $P(X_{1},\ldots,X_{n})$ are implied by so-called d-separations in $G$. A path $X_{i},\ldots,X_{j}$ in $G$ is called blocked by some set $S\subset V\setminus\{X_{i},X_{j}\}$ if every non-collider on the path is in $S$ and every collider and their descendants are not in $S$. If all paths between $X_{i}$ and $X_{j}$ are blocked by $S$ in $G$, we say that $S$ d-separates $X_{i}$ and $X_{j}$ in $G$, denoted by $X_{i}\perp\!\!\!\perp_{G}X_{j}\mid S$. Shortly, $P(X_{1},\ldots,X_{n})$ satisfies the Markov property w.r.t. $G$ if $X_{i}\perp\!\!\!\perp_{G}X_{j}\mid S\Rightarrow X_{i}\perp\!\!\!\perp X_{j}\mid S$ in $P(X_{1},\ldots,X_{n})$.
For a one-to-one correspondence between d-separations in a graph $G=(V,E)$ and conditional independencies in a distribution $P(X_{1},\ldots,X_{n})$ to hold also requires the converse implication—called faithfulness. Faithfulness is not generally satisfied for the observational distribution and graph entailed by an SCM, but often assumed to hold in practice. The subtleties of the faithfulness assumption and its violations have been studied extensively in the causal discovery literature (zhang2002strong; ramsey2012adjacency; uhler2013geometry; marx2021weakerfaithfulnessassumptionbased).
Under the assumptions of causal sufficiency—there are no unobserved common causes of any variables in XX—and faithfulness, d-separations and conditional independencies are in one-to-one correspondence enabling constraint based causal discovery: by testing all possible conditional independencies in $P(X_{1},\ldots,X_{n})$ one can derive all d-separations that hold in $G$. It turns out that the set of DAGs that satisfies a given set of d-separations, called the Markov equivalence class, can be characterized as follows: Each DAG in the Markov equivalence class as the same skeleton (the same set of adjacencies) and the same set of v-structures (pearl2009causal; zhang2008causal). The Markov equivalence class can be represented by a partially directed graph (some directed, some undirected edges). If causal sufficiency is not assumed, the situation becomes more complicated and one can only infer more general types of graphs by testing conditional independencies. The two landmark algorithms for constraint based causal discovery with and without causal sufficiency are the PC and FCI algorithm, respectively. We will first focus on leveraging prior knowledge for a PC type algorithm in the fully observed case, before covering partial observations and an extension of FCI in Section˜3.2.
2 Cluster-DAGs
Cluster-DAGs (C-DAGs) were introduced by anand2023causal as a way to perform causal inference when one can not specify the entire DAG, but has enough information to organize groups of variables into a DAG. C-DAGs aggregate variables into clusters and then define macro-relationships between these clusters. An example can be seen in Fig.˜1.
No assumptions or restrictions are imposed on connections between variables within the same cluster. A directed edge between clusters $C_{1}\rightarrow C_{2}$ means that for any vertices $X\in C_{1},Y\in C_{2}$ there is either no edge between them or it is oriented as $X\rightarrow Y$. If there is no arrow between $C_{1},C_{2}$, no nodes $X\in C_{1},Y\in C_{2}$ are adjacent. Formally, a compatible C-DAG is defined as follows.
Definition 1 (C-DAG, (anand2023causal)).
Given an ADMG $G=(V,E)$ (a graph with directed and bidirected edges as in Definition˜13) and a partition $C={C_{1},\ldots,C_{r}}$ of $V$ (i.e., $C_{i}\cap C_{j}=\emptyset$ for all $i\neq j$ and $V=\bigcup_{i=1}^{r}C_{i}$), construct a graph $G_{C}=(C,E_{C})$ over $C$ with a set of edges $E_{C}$ defined as follows:
  1. (i)
    An edge $C_{i}\rightarrow C_{j}$ is in $E_{C}$ if there exist $X\in C_{i},Y\in C_{j}$ such that $(X\rightarrow Y)\in E$.
  2. (ii)
    A bidirected edge $C_{i}\leftrightarrow C_{j}$ is in $E_{C}$ if there exist $X\in C_{i},Y\in C_{j}$ such that $(X\leftrightarrow Y)\in E$.
If the graph $G_{C}$ contains no directed cycles, $C$ is called an admissible partition of $V$ and $G_{C}$ is called a C-DAG compatible with $G$. Any ADMG $G$ that has $G_{C}$ as a compatible C-DAG is in turn called compatible with $G_{C}$.
$C_{1}$$C_{2}$$C_{3}$
$X_{1}$$X_{3}$$X_{4}$$X_{5}$$X_{2}$$C_{1}$$C_{2}$$C_{3}$
Figure 1: An example of a C-DAG $G_{C}$ and a compatible DAG $G$. Left: A C-DAG $G_{C}$ over three clusters. Right: A graph $G$ compatible with the C-DAG $G_{C}$.
Many different DAGs may be compatible with the same C-DAG and vice versa, so C-DAGs form an equivalence class of DAGs. anand2023causal show that causal effects are still identifiable from the C-DAG $G_{C}$ if they are identifiable in every DAG $G$ compatible with $G_{C}$. Another important result is the soundness and completeness of d-separation in C-DAGs: any d-separation in the C-DAG also holds in any compatible DAG.
Theorem 1 (Soundness and completeness of d-separation in C-DAGs, (anand2023causal)).
In a C-DAG $G_{C}$, let $C_{i},C_{j},C_{k}\subset C$ be sets of clusters. If $C_{i}$ and $C_{j}$ are d-separated by $C_{k}$ in $G_{C}$ (see Definition˜14), then in any ADMG $G$ (see Definition˜13) compatible with $G_{C}$, $C_{i}$ and $C_{j}$ are d-separated by $C_{k}$ in $G$, that is
$C_{i}\perp\!\!\!\perp_{G_{C}}C_{j}\mid C_{k}\quad\Longrightarrow\quad C_{i}\perp\!\!\!\perp_{G}C_{j}\mid C_{k}\>.$ (1)
If $C_{i}$ and $C_{j}$ are not d-separated by $C_{k}$ in $G_{C}$, then there exists an ADMG $G$ compatible with $G_{C}$ where $C_{i}$ and $C_{j}$ are not d-separated by $C_{k}$ in $G$.
Based on these results, anand2023causal then develop an ID algorithm that is sound an complete for identifying causal effects in C-DAGs. In this work, we leverage implications of Theorem˜1 for causal discovery instead.
In constraint-based causal discovery (spirtes2000causation; zhang2008completeness), one typically assumes faithfulness and removes edges from an initially fully connected graph whenever a conditional independence is found between any pair of vertices. The key idea is that Theorem˜1 suggests that whenever $X\in C_{i},Y\in C_{j}$ and $C_{i}\perp\!\!\!\perp_{G_{C}}C_{j}\mid S$, then in any $G$ compatible with $G_{C}$ we have $X\perp\!\!\!\perp_{G}Y\mid S$ (slightly abusing notation in that $S$ is a set of clusters or a union over them, respectively) and thus $X,Y$ cannot be adjacent. Hence, C-DAGs allow us to aggressively prune a fully connected graph to warm-start constraint-based causal discovery on the micro-variables. Besides this warm start that potentially saves many conditional independence tests, the C-DAG structure can also be used to further speed up the remaining discovery process, which we discuss in Section˜3.
2.1 Comparison to tiered background knowledge
Before developing our full causal discovery algorithms with C-DAGs, we provide a thorough comparison with existing forms of groupwise background knowledge—tiered background knowledge (TBK). Like C-DAGs, TBK groups variables, where groups are called tiers. Unlike C-DAGs, tiers impose a directed, chronological ordering between groups of variables.
Definition 2 (Tiered background knowledge, (andrews2020completeness)).
A MAG (maximal ancestral graph, Definitions˜17 and 18) satisfies tiered background knowledge if the variables can be partitioned into $n>1$ disjoint subsets (tiers) $T=\{T_{1},\ldots,T_{m}\}$ and for all $A\in T_{i}$ and $B\in T_{j}$ with $1\leq i<j\leq m$ either (i) $A$ is an ancestor of $B$ or (ii) $A$ and $B$ are not adjacent.
Similar to C-DAGs, TBK enforces the orientations of certain edges due to the macro-constraints. While causal discovery algorithms for TBK like FCI-Tiers (andrews2020completeness) have been developed, C-DAGs are strictly more flexible than TBK:
  1. (i)
    TBK cannot represent settings like the C-DAG $C_{1}\rightarrow C_{3}\leftarrow C_{2}$. TBK would have to put either $C_{1}$ or $C_{2}$ first in the tier ordering, which would only restrict the orientiation of edges between $C_{1},C_{2}$, but it cannot encode the strict absence of edges between $C_{1},C_{2}$ as the C-DAG does.
  2. (ii)
    C-DAGs allow for bidirected edges between clusters, whereas TBK does not, because a bidirected edge could violate both conditions of Definition˜2.
3 Causal discovery with C-DAGs
In this section, we study how C-DAGs with and without unobserved confounding can be incorporated as prior knowledge to improve efficiency and accuracy of constraint-based causal discovery on the micro-variables.
3.1 Cluster-PC
First, we consider C-DAGs without latent confounding, neither within nor between clusters. The assumed C-DAG structure over the micro variables then allows for immediate pruning of the intial complete graph and also for partial orientation of edges via Algorithm˜1. The resulting graph is an MPDAG (see Definition˜11), which can contain both directed and undirected edges. A detailed explanation can be found in Section˜A.1.
1: C-DAG $G_{C}=(V_{C},E_{C})$, $\mathcal{C}=\{C_{1},\ldots,C_{r}\}$
2: Form fully connected graph $G$ over $\cup_{i\in[r]}C_{i}$
3: for $C_{i},C_{j}\in\mathcal{C}$ do
4:   for $X\in C_{i},Y\in C_{j}$ do
5:     if $C_{i}\rightarrow C_{j}$ then orient $X\rightarrow Y$ in $G$.
6:     if $C_{i}\leftarrow C_{j}$ then orient $X\leftarrow Y$ in $G$.
7:     if $C_{i}\cancel{-}C_{j}$ then delete edge XYX-Y in $G$.
8: return MPDAG $G$.
Algorithm 1 C-DAG to MPDAG
Instead of the fully connected graph, this MPDAG will serve as a starting point for our constraint based causal discovery algorithm (akin to PC). The reduced number of adjacencies and partially directed edges reduce the number of required conditional independence (CI) tests during the Cluster-PC algorithm in three ways: (i) directly, as non-adjacent variables are not tested for (conditional) independence anymore, (ii) reducing the number of potential separating sets to be considered, and (iii) by working along a topological ordering of the clusters, more CI tests can be avoided. As an example for the latter, consider the C-DAG $C_{1}\rightarrow C_{2}$. For $X,Y\in C_{1}$ we only need to consider separation sets $S\subset C_{1}$, as any path going through $C_{2}$ contains a collider in $C_{2}$, which is blocked when $C_{2}\cap S=\emptyset$. More generally, the only candidates for a separating set between $X\in C_{1}$ and $Y\in C_{2}$ for the C-DAG $C_{1}\leftarrow C_{2}$ are subsets of the potential parents of $X$ in $C_{1}$ due to Proposition˜1. We now define the general notion of relevant potential separation sets to consider.
Definition 3 (Non-child).
In a PDAG $G$, the non-children of a node $X$ is the set of adjacent nodes $adj_{X}^{G}$ of $X$ that are not children of $X$, i.e., $nch_{X}^{G}:=adj_{X}^{G}\setminus ch_{X}^{G}$.
We can now state the full Cluster-PC (C-PC) algorithm in Algorithm˜2 and show that it is sound and complete.
Theorem 2 (C-PC is sound and complete).
Let $G_{C}$ be a C-DAG compatible with a DAG $G$. Then Algorithm˜2, is sound and complete (for a CI oracle) in that it returns the same MPDAG obtained from running PC on $G$ and orienting all possible additional edges induced by the prior knowledge in $G_{C}$.
The proof can be found in Section˜A.3. When a CI oracle is available, C-PC and PC with post-processing according to the C-DAG return the same result, so C-PC seemingly adds no value. However, in practice CI testing is an inherently difficult problem (ShahPeters2020hardness; Lundborg_2022) and a whole line of works investigates how the number of CI tests can be reduced to render constraint based causal discovery more effective (xie2008recursive; zhang2024membership; shiragur2024causal). In Section˜4 we demonstrate that avoiding unnecessary CI tests by incorporating the prior knowledge in C-PC indeed leads to substantial performance improvements.
Definition 4 (Compatibility of CDPAGs and MPDAGs with C-DAGs).
Let $G=(V,E)$ be the MPDAG obtained from applying Algorithm˜1 to C-DAG $G_{C}$. A CPDAG $G^{\prime}=(V^{\prime},E^{\prime})$ is called compatible with C-DAG $G_{C}$ if and only if $V^{\prime}=V$ and for any $X,Y\in V^{\prime}$ the following holds:
  1. (i)
    If $(X-Y)\in E^{\prime}\Rightarrow(X-Y)\in E$.
  2. (ii)
    If $(X\rightarrow Y)\in E^{\prime}\Rightarrow(X\rightarrow Y)\in E$ (analogous for $(X\leftarrow Y)\in E^{\prime}$).
A graph $G^{\prime}$ being compatible with $G_{C}$ means it can be generated from an algorithm doing edge deletions and orientations on $G$.
Algorithm 2 Cluster-PC algorithm
1: joint distribution $P_{X}$ over $d$ variables Markov and faithful w.r.t. the ground truth graph, CI oracle, C-DAG $G_{C}=(V_{C},E_{C}),C=\{C_{1},\ldots,C_{r}\}$ with clusters in topological ordering and $G_{C}$ compatible with the ground truth CPDAG, see Definition˜4.
2: Get MPDAG $G=(V,E)$ from Algorithm˜1. $\triangleright$ pa, ch, an, de, nb, sib and nch refer to this current $G$
3: for $m\in[r]$ do
4:   $L_{m}\leftarrow C_{m}\cup\bigcup_{C_{s}\in pa_{C_{m}}^{G_{C}}}C_{s}$
5:   for $k=0,\ldots,|L_{m}|-2$ do
6:     $del_{E}\leftarrow\emptyset$
7:     for all pairs $X_{j}\in C_{m}$ and $X_{i}\in pa_{X_{j}}$ do
8:      for all $S\subset nch_{X_{j}}\setminus\{X_{i}\}$ with $|S|=k$ do
9:       if $X_{i}\perp\!\!\!\perp X_{j}\;|\;S$ then
10:          $del_{E}\leftarrow(X_{i}\rightarrow X_{j})$
11:          $S_{\{i,j\}}\leftarrow S$
12:     for all adjacent $X_{i},X_{j}\in C_{m}$ do
13:      for all $S\subset nch_{X_{j}}\setminus\{X_{i}\}$ or $S\subset nch_{i}\setminus\{X_{j}\}$ with $|S|=k$ do
14:       if $X_{i}\perp\!\!\!\perp X_{j}\;|\;S$ then
15:          $del_{E}\leftarrow\{(X_{i}\rightarrow X_{j}),(X_{j}\leftarrow X_{i})$}
16:          $S_{\{i,j\}}\leftarrow S$
17:     $E\leftarrow E\setminus del_{E}$
18: for each triple $X_{i},X_{j},X_{k}\in V$ with $X_{i}-X_{j}-X_{k}$, $X_{i}-X_{j}\leftarrow X_{k}$ or $X_{i}\rightarrow X_{j}-X_{k}$ and $X_{i}\cancel{-}X_{k}$ do $\triangleright$ find v-structures
19:   if $X_{j}\notin S_{\{i,k\}}$ then
20:     orient the edges as $X_{i}\rightarrow X_{j}\leftarrow X_{k}$
21: Successively apply Meek’s edge orientation rules, see Fig.˜5.
22: return MPDAG $G=(V,E)$
3.2 Cluster-FCI
Next, we turn to the partially observed setting, where there may exist latent confounders between observed variables, represented by bidirected edges. Here, we allow bidirected edges both between the actual variables as well as between clusters of variables in the C-DAG. In Section˜3.1, going from the C-DAG to an initial pruned and partially oriented graph was straightforward. This is no longer the case with latent variables, as the cluster graph may now be an ADMG as well (see Definition˜13). We thus refer to such cluster graphs as C-ADMGs. Operating causal discovery directly on ADMGs is a dead end in that one can not distinguish whether there is one or two edges between a pair of nodes. Intuitively, if there may be latent confounders, it is generally impossible to determine whether observed dependence is due to a direct effect or due to latent confounding. In addition, in an ADMG nodes can be non-adjacent while still not being m-separable (m-separation is the natural extension of d-separation to graphs with bidirected edges, see Definition˜16) due to the existence of inducing paths (Definition˜19).
Instead of ADMGs, the literature has converged on ancestral graphs (see Definition˜17) as the core objects in constraint based causal discovery with latent confounders. In our case, we first derive a partial mixed graph from the given C-ADMG and operate our Cluster-FCI (C-FCI) algorithm on it, which ultimately outputs a partial ancestral graph (see, Definition˜21). This output can be viewed as the analogue of the CPDAG in the fully observed setting.
Algorithm 3 C-ADMG to partial mixed graph transformation for C-FCI
1: C-ADMG $G_{C}$ with $C=\{C_{1},\ldots,C_{r}\}$
2: initialize $G$ as a complete undirected graph over $V=\bigcup_{i=1}^{r}C_{i}$
3: for all clusters $C_{i}$ in $G_{C}$ do
4:   for all $X,Y\in C_{i}$ do
5:     add $X\circ\hskip-3.6999pt-\hskip-3.6999pt\circ Y$ to $G$
6: for all adjacent clusters $C_{i},C_{j}$ in $G_{C}$ do
7:   for all $X\in C_{i},Y\in C_{j}$ do
8:     if $C_{i}\rightarrow C_{j}$ and $C_{i}\cancel{\leftrightarrow}C_{j}$ then $E\leftarrow E\cup\{X\rightarrow Y\}$
9:     if $C_{i}\leftarrow C_{j}$ and $C_{i}\cancel{\leftrightarrow}C_{j}$ then $E\leftarrow E\cup\{X\leftarrow Y\}$
10:     if $C_{i}\rightarrow C_{j}$ and $C_{i}\leftrightarrow C_{j}$ then $E\leftarrow E\cup\{X\circ\hskip-4.16243pt\rightarrow Y\}$
11:     if $C_{i}\leftarrow C_{j}$ and $C_{i}\leftrightarrow C_{j}$ then $E\leftarrow E\cup\{X\leftarrow\hskip-4.16243pt\circ Y\}$
12:     if $C_{i}\cancel{\rightarrow}C_{j}$, $C_{i}\cancel{\leftarrow}C_{j}$ and $C_{i}\leftrightarrow C_{j}$ then $E\leftarrow E\cup\{X\leftrightarrow Y\}$
13: for all non-adjacent clusters $C_{i},C_{j}$ connected by an inducing path in $G_{C}$ do
14:   if $C_{i}\in an_{C_{j}}^{G_{C}}$ then
15:     for all $X\in C_{i},Y\in C_{j}$ do $E\leftarrow E\cup\{X\rightarrow Y\}$
16:   if $C_{j}\in an_{C_{i}}^{G_{C}}$ then
17:     for all $X\in C_{i},Y\in C_{j}$ do $E\leftarrow E\cup\{X\leftarrow Y\}$
18:   if $C_{i}\notin an_{C_{j}}^{G_{C}}$ and $C_{j}\notin an_{C_{i}}^{G_{C}}$ then
19:     for all $X\in C_{i},Y\in C_{j}$ do $E\leftarrow E\cup\{X\leftrightarrow Y\}$
20: return partial mixed graph $G_{pm}:=G$
Definition 5 (Partial mixed graph for C-ADMG).
[…]
  1. (i)
    for all $X,Y\in C_{i}$ we have $(X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y)\in E$,
  2. (ii)
    for all $X\in C_{i},Y\in C_{j}$ with $C_{i}\rightarrow C_{j}$ and $C_{i}\not\leftrightarrow C_{j}$ we have $(X\rightarrow Y)\in E$,
  3. (iii)
    for all $X\in C_{i},Y\in C_{j}$ with $C_{i}\rightarrow C_{j}$ and $C_{i}\leftrightarrow C_{j}$ we have $(X\circ\hskip-4.49997pt\rightarrow Y)\in E$,
  4. (iv)
    for $X\in C_{i},Y\in C_{j}$ with $C_{i},C_{j}$ not adjacent and connected by an inducing path, if
    • $C_{i}\in an_{C_{j}}^{G_{C}}$ we have $(X\rightarrow Y)\in E$,
    • $C_{j}\in an_{C_{i}}^{G_{C}}$ it is $(X\leftarrow Y)\in E$,
    • $C_{i}\notin an_{C_{j}}^{G_{C}}$ and $C_{j}\notin an_{C_{i}}^{G_{C}}$ we have $(X\leftrightarrow Y)\in E$.
The circle indicates uncertainty on the nature of the edge mark; it could be a tail or an arrow. In MPDAGs, an undirected edge XYX-Y was used to express uncertainty about the edge direction. In partial mixed graphs this function is now served by the edge $X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y$.
For brevity of notation, as in zhang2008completeness, we also introduce the "*" sympol to denote either an arrowhead, circle or tail. This will be used later for edge orientations, where some edge marks don’t matter for the orientation rule. For example if the edge $X*\hskip-5.0pt-\hskip-3.99994pt\circ Y$ is oriented as $X*\hskip-5.0pt\rightarrow Y$, the edge mark on the left stays whatever it was before the orientation. In the original definition of partial ancestral graphs (zhang2008completeness), undirected edges "-" or $\circ\hskip-1.79993pt-"$ are present in order to consider potential selection variables. In our work we assume non-existence of selection variables, so we omit these undirected edges here.
The partial mixed graph $G_{pm}$ for a C-ADMG $G_{C}$ will be the starting graph for the Cluster-FCI Algorithm˜4 and is produced from $G_{C}$ by using Algorithm˜3. To obtain an ancestral graph later on during C-FCI, non-adjacent but non-m-separable nodes need to have an edge introduced between them in the preprocessing, see Algorithm˜3(l. 12-18).
$X_{1}$$X_{2}$$X_{3}$$X_{4}$$X_{5}$$C_{1}$$C_{2}$$C_{3}$
Figure 2: Example of non-ancestral C-ADMG.
Algorithm 4 Cluster-FCI algorithm
1: Joint distribution $P_{O}$ of $d$ observed variables, independence oracle, C-DAG $G_{C}=(V_{C},E_{C}),C=\{C_{1},\ldots,C_{r}\}$ with clusters in topological ordering (w.r.t. directed edges) compatible with ground truth MAG, see Definition˜6.
2: Construct graph $G=(V,E)$ from $G_{C}$ with Algorithm˜3.
3: for $m\in[r]$ do
4:   $L_{m}\leftarrow C_{m}\cup\bigcup_{C_{s}\in pa_{C_{m}}^{G_{C}}}C_{s}\cup\bigcup_{C_{s}\in sib_{C_{m}}^{G_{C}}}C_{s}$
5:   for $k=0,\ldots,|L_{m}|-2$ do
6:     $del_{e}\leftarrow\emptyset$
7:     for for all $X_{i}\in C_{m}$ and $X_{j}\in nch_{X_{i}}$ do
8:      for all $S\subset nch_{X_{j}}\setminus\{X_{i}\}$ or $S\subset nch_{X_{i}}\setminus\{X_{j}\}$ with $|S|=k$ do
9:       if $X_{i}\perp\!\!\!\perp X_{j}\mid S$ then
10:          $del_{E}\leftarrow del_{E}\cup\{X_{i}*\hskip-4.62497pt-\hskip-4.62497pt*X_{j}\}$
11:          $S_{\{i,j\}}\leftarrow S$
12:     $V\leftarrow V\setminus del_{E}$
13: for all unshielded triples $(X_{i},X_{j},X_{k})$ do
14:   if $X_{j}\notin S_{\{i,k\}}$ then
15:     if $X_{i}*\hskip-4.62497pt-\hskip-3.6999pt\circ X_{j}\circ\hskip-3.6999pt-\hskip-4.62497pt*X_{k}$ then orient $X_{i}*\hskip-4.62497pt-\hskip-3.6999pt\circ X_{j}\circ\hskip-3.6999pt-\hskip-4.62497pt*X_{k}$ as $X_{i}*\hskip-4.62497pt\rightarrow X_{j}\leftarrow\hskip-4.62497pt*X_{k}$ in $G$
16:     if $X_{i}*\hskip-4.62497pt\rightarrow X_{j}\circ\hskip-4.16243pt\rightarrow X_{k}$ then orient $X_{i}*\hskip-4.62497pt\rightarrow X_{j}\circ\hskip-4.16243pt\rightarrow X_{k}$ as $X_{i}*\hskip-4.62497pt\rightarrow X_{j}\leftrightarrow X_{k}$ in $G$
17: for all $X_{i}\in V$ do
18:   for all $X_{j}\in adj_{X_{i}}^{G}$ do
19:     compute $pds(X_{i},X_{j})$ as in Definition˜22.
20:     for $k=0,\ldots,d-2$ do
21:      for $|S|\subset pds(X_{i},X_{j})$ with $|S|=k$ do
22:       if $X_{i}\perp\!\!\!\perp X_{j}\mid S$ then
23:          $V\leftarrow V\setminus\{X_{i}*\hskip-4.62497pt-\hskip-4.62497pt*X_{j}\}$
24:          $S_{\{i,j\}}\leftarrow S$
25: Reorient all edges according to C-DAG $G_{C}$ (as in Algorithm˜3 but only orienting edges, not adding edges)
26: For any almost directed cycle $X_{l}\leftrightarrow X_{i}\rightarrow\ldots\rightarrow X_{l}$ orient $X_{l}\leftrightarrow X_{i}$ to $X_{l}\rightarrow X_{i}$
27: Use rules R0-R4, R8-R10 of (zhang2008completeness) to orient as many edge marks as possible.
28: return PAG $G=(V,E)$
Definition 6 (Compatibility of MAGs and partial mixed graphs with C-ADMGs).
Let $G_{pm}=(V,E)$ be the partial mixed graph obtained from applying Algorithm˜3 to C-ADMG $G_{C}$. A partial mixed graph $G_{pm}^{\prime}=(V^{\prime},E^{\prime})$ is called compatible with C-ADMG $G_{C}$ if $V^{\prime}=V$ and for any $X,Y\in V^{\prime}$ the following holds:
  1. (i)
    $(X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y)\in E^{\prime}\Rightarrow(X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y)\in E$,
  2. (ii)
    $(X\circ\hskip-4.49997pt\rightarrow Y)\in E^{\prime}\Rightarrow\{(X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y),(X\circ\hskip-4.49997pt\rightarrow Y)\}\cap E\neq\emptyset$ (analogous for $(X\leftarrow\hskip-4.49997pt\circ Y)\in E^{\prime}$),
  3. (iii)
    $(X\leftrightarrow Y)\in E^{\prime}\Rightarrow\{(X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y),(X\circ\hskip-4.49997pt\rightarrow Y),(X\leftarrow\hskip-4.49997pt\circ Y),(X\leftrightarrow Y)\}\cap E\neq\emptyset$,
  4. (iv)
    $(X\rightarrow Y)\in E^{\prime}\Rightarrow X\in nch_{Y}^{G_{pm}}$ (analogous for $(X\leftarrow Y)\in E^{\prime}$).
Compatibility of MAGs (which are partial mixed graphs that satisfy ancestrality and maximality (zhang2008completeness)) with C-ADMGs follows directly from this definition, too. To summarize, a graph $G_{pm}^{\prime}$ being compatible with $G_{C}$ means it can be generated from an algorithm doing edge deletions and orientations on $G_{pm}$.
The graph resulting from Algorithm˜3 need not necessarily be ancestral yet, as the input C-ADMG, and thus also the output, may contain almost directed cycles. An almost directed cycle is defined as $X\leftrightarrow Y$ and $X\in an_{Y}^{G}$ (zhang2008completeness). A direct conversion of the C-ADMG to a MAG, e.g., following hu2020fasteralgorithmsmarkovequivalence, is undesirable, as the following example demonstrates. Consider the graph in Fig.˜2. The MAG $G_{M}$ over $\{X_{1},\ldots,X_{5}\}$ is ancestral, but the corresponding C-ADMG $G_{C}$ with $C=\{C_{1},C_{2},C_{3}\},C_{1}=\{X_{1},X_{2}\},C_{2}=\{X_{3}\},C_{3}=\{X_{4},X_{5}\}$ is not, due to the almost directed cycle $C_{3}\leftrightarrow C_{1}\rightarrow C_{2}\rightarrow C_{3}$. Even though there is an almost directed cycle in $G_{C}$, $G_{M}$ is compatible with $G_{C}$. If we were to transform this C-ADMG into a MAG, the edge $C_{1}\leftrightarrow C_{2}$ would change to $C_{1}\rightarrow C_{2}$. The edge between $X_{2},X_{5}$ would consequently inherit this orientation as $X_{2}\rightarrow X_{5}$ contradicting the correct edge type $X_{2}\leftrightarrow X_{5}$. Therefore, we only reorient almost directed cycles at the very end of C-FCI, ensuring it outputs a valid PAG without introducing false edge information.
Definition 7 (Updated pa, ch and nch, and sib).
In addition to the previous definitions, in the following, whenever $X\circ\hskip-4.49997pt\rightarrow Y$, we count $X$ as a parent of $Y$, $X\in pa_{Y}$, and vice versa $Y\in ch_{X}$. Whenever $X\circ\hskip-3.99994pt-\hskip-3.99994pt\circ Y$, $X\circ\hskip-4.49997pt\rightarrow Y$, or $X\leftrightarrow Y$, we count $X$ as a non-child of $Y$, $X\in nch_{Y}$ and these definitions extend to ancestors and descendants. Whenever $X\leftrightarrow Y$, we call $X$ is a sibling of $Y$, $X\in sib_{Y}$ (and vice versa).
Cluster-FCI follows a similar strategy as Cluster-PC by running CI tests on a per-cluster basis (while following the general logic of FCI) and is developed in all detail in Algorithm˜4.
We now highlight some differences between Cluster-FCI and FCITiers. Cluster-FCI starts with the first cluster of the topological ordering and works its way down along the topological ordering, while FCITiers (andrews2020completeness, Alg. 1) starts with the last and works its way up. However, due to the nature of FCITiers, this direction does not matter as it is running a version of FCI on disjoint sets of edges (FCIExogenous). These disjoint edge sets are derived from the TBK. The resulting edges from FCIExogenous are then added to the overall graph. On the contrary, Cluster-FCI pre-processes the entire fully connected graph according to the C-ADMG to obtain a partial mixed graph. It then removes further edges along the topological ordering. Our proposed C-FCI Algorithm˜4 is sound, but not complete (proof in Section˜A.4).
Theorem 3 (Soundness of Cluster-FCI).
If the C-DAG $G_{C}$ is compatible with ground truth MAG $G_{M}$, C-FCI is sound in the sense that nodes $X_{i},X_{j}$ are adjacent in the output PAG $G_{P}$ if and only if they are adjacent in the ground truth MAG $G_{M}$. In addition, all arrow and tail edge marks in $G_{P}$ are also present in $G_{M}$.
Too much causal information—incompleteness of C-FCI.  The example in Fig.˜2 shows that a C-ADMG can encode arrowheads that contradict ancestrality (imagine an additional edge $X_{1}\leftrightarrow X_{4}$). C-FCI can be adjusted to not output a PAG (Non-PAG C-FCI), by not re-orienting almost directed cycles. This variation can be an improvement, as it increases the information contained in the obtained graph. To see this, consider the graph in Fig.˜2 with an additional edge $X_{1}\leftrightarrow X_{4}$. C-FCI would return the edge $X_{1}\rightarrow X_{4}$, due to the almost directed cycle $X_{4}\leftrightarrow X_{1}\rightarrow X_{3}\rightarrow X_{4}$. Non-PAG C-FCI in contrast will return $X_{1}\leftrightarrow X_{4}$, the correct and more informative result. Further exploring this deviation from the well-explored setting of relying primarily on ancestral graphs for causal discovery is an interesting direction for future work.
This example simultaneously highlights that C-FCI is incomplete: its output does not always reveal all determined causal information. In a way, C-FCI is “overwhelmed by prior causal information,” as some (useful) background knowledge can violate ancestrality and thus not be captured properly in the algorithm. However, C-FCI always remains at least as informative as FCI. In Remark˜1 we additionally sketch that C-FCI is also at least as informative as FCITiers for a suitable set of tiers—recalling that C-DAGs are strictly more flexible then TBK. Finally, we conjecture that for ancestral C-ADMGs, C-FCI is also complete, i.e., all counterexamples have to rely on non-ancestral C-ADMGs as background knowledge. This also remains an open question for future work.
Figure 3: Left: Precision vs. recall for different significance levels $\alpha$ of the CI test. Cluster-PC dominates base PC w.r.t. recall and for common values of $\alpha\in\{0.05,0.01\}$ achieves substantially better recall with only small reductions in precision. For details, see Tables˜4 and 5. Right: Precision, recall and F1-score for different numbers of clusters. The performance gap grows with the number of clusters, which amounts to more granular background knowledge.
Figure 4: Left: Structural Hamming distance (SHD) for different numbers of clusters. The empty graph is used as a dummy reference. Again, C-PC clearly benefits from a growing number of clusters. Middle: The number of conditional independence (CI) tests is substantially reduced in C-PC for C-DAGs with many clusters. Notably, even for coarse background knowledge of only two clusters, the number of required CI tests already drops noticeably. Right: The ratio of CI tests between PC and C-PC (for the same graph, across different numbers of edges and clusters), highlights that the savings remain roughly constant for a fixed number of clusters even as the number of edges increases.
4 Simulation studies
We now empirically demonstrate the differences between PC vs. Cluster-PC and FCI vs. FCITiers vs. Cluster-FCI in different simulation studies. All code for these experiments is available at https://github.com/JanMarcoRuizdeVargas/clustercausal/tree/main . In the first setting, we sample 1750 Erdős–Rényi graphs and vary the number of nodes, edges, and the significance level $\alpha$ for the chosen CI tests. The second simulation study performs sensitivity analysis w.r.t. different graph generation methods and probability distributions. The third compares the three FCI variants using generated C-ADMGs (that not necessarily satisfy TBK). The last simulation study compares the FCI variants on C-DAGs, which do satisfy TBK. Section˜B.2 contains a detailed breakdown of all chosen simulation parameters and settings.
We evaluate the discovery algorithms with respect to precision, recall, F1-score, Structural Hamming Distance (SHD), and the number of required CI-tests. For these metrics, we distinguish between ‘adjacency’, i.e., is there any edge present between two variables, and ‘arrow’, which compares the types of edge marks. Detailed definitions of the used metrics can be found in Appendix˜B. We also run the algorithms with different significance levels $\alpha$. Since rejecting the null hypothesis, i.e., rejecting conditional independence, leads to an edge not being removed, while failure to reject leads to a removed edge, higher $\alpha$ leads to fewer edge deletions and overall denser graphs. On the contrary, smaller $\alpha$ leads to more deletions and thus sparser final graphs.
Table 1: Comparison of metrics for Simulations 1 (left) and 2 (right). C-PC considerably outperforms PC in the arrow metrics. The cluster version is slightly worse in adjacency precision, but shows improvements in adjacency recall, F1-score, SHD and reduces the number of CI tests required.
Metric PC Cluster-PC
Adj. precision 87.9% 85.1%
Arrow precision 54.4% 71.8%
Adj. recall 46.5% 55.5%
Arrow recall 24.2% 45.6%
Adj. F1-score 55.3% 62.7%
Arrow F1-score 30.9% 52.1%
SHD 44.8 38.2
Avg. CI tests 4981 3062
Metric PC Cluster-PC
Adj. precision 86.3% 83.6%
Arrow precision 61.1% 73.1%
Adj. recall 46.4 % 54.5%
Arrow recall 26.4 % 45.2%
Adj. F1-score 55.8% 62.1%
Arrow F1-score 34.4% 52.6%
SHD 40.4 35.4
Avg. CI tests 4580 2995
Table 2: Comparison of metrics between FCI, Cluster-FCI, and FCITiers for Simulations 3 (left) and 4 (right). Simulation 3 generated C-ADMGs, which not necessarily satisfy TBK, while Simulation 4 generated C-DAGs with latent variables only within clusters. This means a topologically ordered, fully connected alteration of such a C-DAG satisfies TBK. There is a difference in the difficulty of causal discovery between the two simulations, as we needed to use a different graph generation method for Simulation 4 (see Section˜B.3). So the metrics should only be compared within a simulation study, not across. On C-ADMGs, C-FCI outperforms FCI and FCITiers, while being close in performance to FCITiers on C-DAGs with latent variables within clusters.
Metric FCI C-FCI FCITiers
Adj. precision 36.0% 35.2% 32.5%
Arrow precision 24.5% 29.9% 29.8%
Adj. recall 21.5% 24.8% 22.8%
Arrow recall 11.9% 20.1% 16.7%
Adj. F1-score 26.3% 28.6% 26.3%
Arrow F1-score 15.3% 23.6% 20.9%
SHD 29.3 30.4 29.89
CI tests 1008 559 1139
Metric FCI C-FCI FCITiers
Adj. precision 94.9% 93.3% 93.5%
Arrow precision 67.0% 81.1% 81.2%
Adj. recall 62.8% 63.8% 63.8%
Arrow recall 41.0% 43.0% 44.4%
Adj. F1-score 74.8% 75.0% 75.1%
Arrow F1-score 49.5% 55.1% 56.4%
SHD 20.2 16.1 16.6
Avg. CI tests 1475 823 1568
Figs.˜3 and 1 demonstrate that C-PC dominates PC in recall as well as arrow precision and F1-score. A mild disadvantage in adjacency precision is minor compared to the gains in recall and can mostly be removed by choosing smaller $\alpha$. Fig.˜4 further shows that the efficiency, i.e., the number of CI tests, drastically improved with more fine grained background knowledge (more clusters), while also improving the overall performance of the causal discovery measured by SHD. Even coarse background knowledge from a C-DAG consisting of just two clusters substantially reduces the number of required CI tests. Finally, the efficiency gains are not sensitive to the overall number of edges in a graph, but only depend on the number of clusters, i.e., the ‘amount of the background information.’
For C-FCI, Table˜2(left) also shows improved accuracy and arrow precision with minor hits in adjacency precision compared to FCI and also to FCITiers in the setting where C-DAGs do not necessarily satisfy TBK. While overall SHD remains comparable, C-FCI requires around half the number of CI tests. In the setting, where generated C-DAGs only have latent variables within clusteres, i.e., TBK can be satisfied according to some topologically ordered completion of the C-DAG (Table˜2(right)), C-FCI and FCITiers perform similarly (both generally outperforming vanilla FCI) whereas again C-FCI requires only about half the number of CI tests.
5 Conclusion
We leverage C-DAGs as a flexible and realistic type of background knowledge for constraint-based causal discovery in fully and partially observed settings. C-DAGs are a provably superior alternative for encoding group-wise background knowledge compared to the existing tiered background knowledge. We develop the Cluster-PC (C-PC) and Cluster-FCI (C-FCI) algorithms and prove that they are complete and sound or sound (but not complete) respectively. The non-completeness of C-FCI is shown to stem from C-ADMGs possibly containing ‘more causal information’ than the well-established PAG representation in causal discovery with latents can handle. Through extensive empirical simulation studies we demonstrate that our proposed algorithms indeed outperform the corresponding algorithms with no, or existing types of background knowledge across a wide range of settings on most metrics.
Interesting future directions for future work include to apply C-DAG background knowledge to score-based methods (see Appendix˜D for some first thoughts), to combine C-DAGs with other types of prior knowledge such as pairwise background knowledge, or to include selection variables in the causal discover process as well. Since C-ADMGs can contain background knowledge that violates ancestrality required for FCI, it will also be interesting to investigate under which types of background knowledge an extended FCI version remains complete. Lastly, using summary causal graphs (ferreira2025identifyingC)—also allowing for cycles—instead of C-DAGs for improved causal discovery is also an interesting direction for follow up work.
Appendix A Definitions, theorems, additional explanations
A.1 A C-DAG leads to an MPDAG
The Markov equivalence class of a graph $G$, the graphs entailing the same d-separations, can be characterized by a CPDAG.
Definition 8 (CPDAG (completed partially directed graph), andersson1997characterization).
The completed partially directed graph of a DAG $G$, denoted by $G^{*}$, is a graph with the same skeleton as $G$ and undirected edges. A directed edge occurs if and only if that directed edge is present in all DAGs of the Markov equivalence class of $G$. Directed edges come from either v-structures or applying the orientation rules in Fig.˜5.
Figure 5: Meek’s orientation rules (meek2013causalinferencecausalexplanation) (Figure from (fang2022representationcausalbackgroundknowledge)).
Definition 9 (Pairwise Causal Constraints, fang2022representationcausalbackgroundknowledge).
A direct causal constraint, denoted by $X\rightarrow Y$, is a proposition stating that $X$ is a parent of $Y$, i.e., $X$ is a direct cause of $Y$. An ancestral causal constraint, denoted by $X\dashrightarrow Y$, is a proposition stating that $X$ is an ancestor of $Y$, i.e., $X$ is a cause of $Y$. A non-ancestral causal constraint, denoted by $X\cancel{\dashrightarrow}Y$, is a proposition stating that $X$ is not an ancestor of $Y$, i.e., $X$ is not a cause of $Y$. In all these cases, $X$ is called the tail and $Y$ is called the head.
Definition 10 (Restricted Markov equivalence class,fang2022representationcausalbackgroundknowledge).
The restricted Markov equivalence class induced by a CPDAG $G^{*}$ and a pairwise causal constraint set $B$ over $V$, denoted by $G^{*},B]$ is composed of all equivalent DAGs in $M(G^{*})$ that satisfy $B$ ($M(G^{*})$ is the Markov equivalence class of $G^{*}$).
Definition 11 (Maximally partially directed acyclic graph (MPDAG), fang2022representationcausalbackgroundknowledge).
The MPDAG $H$ of a non-empty restricted Markov equivalence class $G^{*},B]$, induced by a CPDAG $G^{*}$ and a pairwise causal constraint set $B$ is a PDAG such that
  1. (i)
    $H$ has the same skeleton and v-structures as $G^{*}$ and
  2. (ii)
    an edge is directed in $H$ if and only if it appears in all DAGs in $G^{*},B]$.
So for a given CPDAG, the C-DAG pairwise constraint set would be (also see Appendix˜C):
Definition 12 (Pairwise causal constraint set from C-DAGs).
The pairwise causal constraint $B$ set induced by C-DAG $G_{C}$ is as follows: For $X_{i}\in C_{i},X_{j},\in C_{j}$, $C_{i}\neq C_{j}$,
  • if $C_{i}\rightarrow C_{j}$, then $X_{i}\cancel{\dashleftarrow}X_{j}\in B$
  • if $C_{i}\dashrightarrow C_{j},C_{i}\cancel{\rightarrow}C_{j}$, then $X_{i}\cancel{\dashleftarrow}X_{j},X_{i}\cancel{\leftarrow}X_{j}\in B$
  • if $C_{i}\cancel{\dashleftarrow}C_{j}$, $C_{i}\cancel{\dashrightarrow}C_{j}$, then $X_{i}\cancel{\dashleftarrow}X_{j},X_{i}\cancel{\dashrightarrow}X_{j}\in B$
C-DAG leads to MPDAG. It follows from the definitions that a CPDAG restricted by a compatible C-DAG leads to an MPDAG. In the same way, restricting a fully connected graph with a C-DAG via Algorithm˜1 also leads to an MPDAG, whose edges are a super-set of any compatible DAG or CPDAG (interpreting undirected edges as $\rightarrow$ and $\leftarrow$).
A.2 Further definitions, theorems and additional explanations
Definition 13 (ADMG).
A directed mixed graph $G=(V,E)$ consists of a finite set of nodes $V$ and a finite set of edges $E$, which are either directed ($\rightarrow$) or bidirected ($\leftrightarrow$). An acyclic directed mixed graph (ADMG) is a directed mixed graph without directed cycles.
Definition 14 (d-separation in C-DAGs,anand2023causal).
A path $p$ in a C-DAG $G_{C}$ is said to be d-separated by a set of clusters $Z\subset C$ if and only if $p$ contains a triplet
  1. (i)
    $C_{i}*\hskip-5.0pt-\hskip-5.0pt*C_{m}\rightarrow C_{j}$ such that the non-collider cluster $C_{m}$ is in $Z$, or
  2. (ii)
    $C_{i}*\hskip-5.0pt\rightarrow C_{m}\leftarrow\hskip-5.0pt*C_{j}$ such that the collider cluster $C_{m}$ and its descendants are not in $Z$.
A set of clusters $Z$ is said to d-separate two sets of clusters $X,Y\subset C$, denoted by $X\perp\!\!\!\perp_{G_{C}}Y\mid Z$, if and only if $Z$ blocks every path from a cluster in $X$ to a cluster in $Y$.
Definition 15 (mns (minimal neighbor separator), gupta2023local).
For a DAG $G=(V,E)$ and node $X$ and $A\notin nb^{+}_{X}$ ($nb^{+}_{X}:=nb_{X}\cup\{X\}$), the minimal neighbor separator $mns_{X}(A)\subset nb_{X}$ is the unique set of nodes such that
  1. (i)
    (d-separation) $A\perp\!\!\!\perp_{G}X\mid mns_{X}(A)$
  2. (ii)
    (minimality) for any $S\subset mns_{X}(A):A\not\!\perp\!\!\!\perp_{G}X\mid S$
hold.
Proposition 1 (Restricting separating set via mns,gupta2023local).
For any node $Y\notin de_{X}\cup nb_{X}^{+}$, the minimum neighbor separator $mns_{X}(Y)$ exists and $mns_{X}(Y)\subset pa_{X}$.
Definition 16 (M-connecting, m-separation).
Let $G=(V,E)$ be a directed mixed graph (a graph containing directed and bidirected edges). A path between $X_{i},X_{j}\in V$ is called m-connecting in $G$ given $S\subset V$ if every non-collider on the path is not in $S$, and every collider on the path is in $S$ or is an ancestor of $S$ in $G$. If there is no path m-connecting $X_{i}$ and $X_{j}$ in $G$ given $S$, $X_{i}$ and $X_{j}$ are called m-separated given $S$. Sets $A$ and $B$ are said to be m-separated given $S$, if for all $X_{i}\in A$ and all $j\in B$, $X_{i}$ and $X_{j}$ are m-separated given $S$.
Definition 17 (Ancestral graph).
A mixed graph $G$ (containing directed and bidirected edges) is ancestral if the following three conditions hold:
  1. (i)
    there is no directed cycle,
  2. (ii)
    there is no almost directed cycle,
  3. (iii)
    for any undirected edge $X_{1}-X_{2}$, $X_{1}$ and $X_{2}$ have no parents or siblings.
Definition 18 (MAG (maximal ancestral graph).
An ancestral graph is called maximal if for any two non-adjacent vertices, there is a set of vertices that m-separates them.
Definition 19 (Inducing path,zhang2008completeness).
In an ancestral graph, let $X,Y$ be any two vertices and $L,S$ be disjoint sets of vertices not containing $X,Y$. $L$ describes the latent variables and $S$ describes the selection variables. A path $\pi$ between $X$ and $Y$ is called an inducing path relative to $(L,S)$ if every non-endpoint vertex on $\pi$ is either in $L$ or a collider, and every collider on $\pi$ is an ancestor of either $X,Y$, or a member of $S$. When $L=S=\emptyset$, $\pi$ is called a primitive inducing path between $X$ and $Y$.
Two DAGs are Markov equivalent if and only if they have the same adjacencies and unshielded colliders. For two MAGs, this is a necessary condition, but not sufficient anymore. For two MAGs to be Markov equivalent, they also need to possess the same colliders on discriminating paths.
Definition 20 (Discriminating path,zhang2008completeness).
In a MAG, a path between $X$ and $Y$, $\pi=(X,\ldots,W,S,Y)$ is a discriminating path for $S$ if
  1. (i)
    $\pi$ includes at least three edges,
  2. (ii)
    $S$ is a non-endpoint vertex on $\pi$ and is adjacent to $Y$ on $\pi$,
  3. (iii)
    $X$ is not adjacent to $Y$ and every vertex between $X$ and $S$ is a collider on $\pi$ and is a parent of $Y$.
Proposition 2 (Markov equivalence criterion for MAGs,zhang2008completeness).
Two MAGs over the same set of vertices are Markov equivalent if and only if
  1. (i)
    they have the same adjacencies,
  2. (ii)
    they have the same unshielded colliders,
  3. (iii)
    if a path $\pi$ is a discriminating path for a vertex $S$ in both graphs, then $S$ is a collider on the path in one graph if and only if it is a collider on the path in the other.
The partial ancestral graph is the CPDAG analogue for characterizing the Markov equivalence class:
Definition 21 (Partial ancestral graph, zhang2008completeness).
[…]
  1. (i)
    $G_{P}$ has the same adjacencies as $G$ (and any member of $M(G)$ and
  2. (ii)
    every non-circle mark in $G_{P}$ is an invariant mark in $M(G)$.
If furthermore every circle in $G_{P}$ corresponds to a variant mark in $M(G)$, $G_{P}$ is called the maximally informative PAG for $M(G)$.
In ancestral graphs, it may be possible that two m-separable nodes can not be m-separated by (a subset of) their neighbors. So FCI needs to search ’possible d-separating sets’ too, i.e., sets that contain nodes not adjacent to $X,Y$, but whose nodes may be necessary to m-separate $X$ and $Y$.
Definition 22 (Possible d-separating set, andrews2020completeness).
$X\in pds(X_{i},X_{j})$ if and only if $X\notin\{X_{i},X_{j}\}$ and there is a path $\pi$ between $X_{i}$ and $X$ in $G$ such that for every subpath $(X_{k},X_{l},X_{m})$ of $\pi$ either $X_{l}$ is a collider on $\pi$ or $X_{k}$ and $X_{m}$ are adjacent.
Definition 23 (Orientation rules for FCI, zhang2008completeness).
  • R0: For each unshielded triple $\langle\alpha,\gamma,\beta\rangle$ in $P$, orient as a collider $\alpha*\hskip-5.0pt\rightarrow\gamma\leftarrow\hskip-5.0pt*\beta$ if and only if $\gamma\notin\text{Sepset}(\alpha,\beta)$.
  • R1: If $\alpha*\hskip-5.0pt\rightarrow\beta\circ\hskip-3.99994pt-\hskip-5.0pt*\gamma$, and $\alpha$ and $\gamma$ are not adjacent, then orient the triple as $\alpha*\hskip-5.0pt\rightarrow\beta\rightarrow\gamma$.
  • R2: If $\alpha\rightarrow\beta*\hskip-5.0pt\rightarrow\gamma$ or $\alpha*\hskip-5.0pt\rightarrow\beta\rightarrow\gamma$, and $\alpha\circ\hskip-3.99994pt-\hskip-3.99994pt\circ\gamma$, then orient $\alpha*\hskip-5.0pt-\hskip-3.99994pt\circ\gamma$ as $\alpha*\hskip-5.0pt\rightarrow\gamma$.
  • R3: If $\alpha*\hskip-5.0pt\rightarrow\beta\leftarrow\hskip-5.0pt*\gamma$, $\alpha*\hskip-5.0pt-\hskip-3.99994pt\circ\theta\circ\hskip-3.99994pt-\hskip-5.0pt*\gamma$, $\alpha$ and $\gamma$ are not adjacent, and $\theta*\hskip-5.0pt-\hskip-3.99994pt\circ\beta$, then orient $\theta*\hskip-5.0pt-\hskip-3.99994pt\circ\beta$ as $\theta*\hskip-5.0pt\rightarrow\beta$.
  • R4: If $u=\langle\theta,\dots,\alpha,\beta,\gamma\rangle$ is a discriminating path between $\theta$ and $\gamma$ for $\beta$, and $\beta\circ\hskip-3.99994pt-\hskip-5.0pt*\gamma$; then if $\beta\in\text{Sepset}(\theta,\gamma)$, orient $\beta\circ\hskip-3.99994pt-\hskip-5.0pt*\gamma$ as $\beta\rightarrow\gamma$; otherwise orient the triple $\langle\alpha,\beta,\gamma\rangle$ as $\alpha\leftrightarrow\beta\leftrightarrow\gamma$.
  • R8: If $\alpha\rightarrow\beta\rightarrow\gamma$ or $\alpha-\hskip-1.79993pt\circ\beta\rightarrow\gamma$, and $\alpha\circ\hskip-4.49997pt\rightarrow\gamma$, orient $\alpha\circ\hskip-4.49997pt\rightarrow\gamma$ as $\alpha\rightarrow\gamma$.
  • R9: If $\alpha\circ\hskip-4.49997pt\rightarrow\gamma$, and $p=\langle\alpha,\beta,\theta,\dots,\gamma\rangle$ is an uncovered p.d. (partially directed) path from $\alpha$ to $\gamma$ such that $\gamma$ and $\beta$ are not adjacent, then orient $\alpha\circ\hskip-4.49997pt\rightarrow\gamma$ as $\alpha\rightarrow\gamma$.
  • R10: Suppose $\alpha\circ\hskip-4.49997pt\rightarrow\gamma$, $\beta\rightarrow\gamma\leftarrow\theta$, $p_{1}$ is an uncovered p.d. path from $\alpha$ to $\beta$, and $p_{2}$ is an uncovered p.d. path from $\alpha$ to $\theta$. Let $\mu$ be the vertex adjacent to $\alpha$ on $p_{1}$ ($\mu$ could be $\beta$), and $\omega$ be the vertex adjacent to $\alpha$ on $p_{2}$ ($\omega$ could be $\theta$). If $\mu$ and $\omega$ are distinct, and are not adjacent, then orient $\alpha\circ\hskip-4.49997pt\rightarrow\gamma$ as $\alpha\rightarrow\gamma$.
A.3 Soundness and completeness of Cluster-PC
Theorem 4 (Soundness and completeness of C-PC).
When the C-DAG $G_{C}$ is compatible with the ground truth DAG $G$, the C-PC algorithm as stated in Algorithm˜2 is sound and complete (when using a CI oracle). Sound and complete in the sense that it returns the same MPDAG as using PC on $G$ and orienting additional edges according to $B$ (the pairwise causal constraint set induced by $G_{C}$).
Proof: Let $\hat{G}$ be the output of the C-PC algorithm and $\bar{G}$ the MPDAG of the restricted Markov equivalence class $G^{*},B]$, obtained by restricting the CPDAG output from PC with $B$. $G$ is the ground truth DAG. $V$ is the node set and $E_{\hat{G}},E_{\bar{G}},E_{G^{*}}$ their edge sets, respectively. First, we show that $\hat{G}$ has the same skeleton as $\bar{G}$. Let $X-Y\in E_{\bar{G}}$ be any edge in $\bar{G}$. This means $X,Y$ are not d-separable in $G$, and thus any compatible C-DAG $G_{C}$ will not put $X,Y$ into non-adjacent clusters. Also, no CI test performed in C-PC will remove this edge by the global Markov property ($\forall S:X\perp\!\!\!\perp_{G}Y|S\Rightarrow X\perp\!\!\!\perp Y|S$, thus $\forall S:X\not\!\perp\!\!\!\perp Y|S\rightarrow X\not\!\perp\!\!\!\perp_{G}Y|S$). On the other hand, let $X,Y$ be non-adjacent in $\bar{G}$, so $\exists S:X\perp\!\!\!\perp Y|S$. W.l.o.g. let $Y\notin de_{X}\cup nb_{X}^{+}$, then by Proposition˜1 $mns_{X}(Y)\subset pa_{X}$ and $X\perp\!\!\!\perp Y|mns_{X}(Y)$. Furthermore, $mns_{X}(Y)\subset pa_{X}\subset nch_{X}$ and for every $S^{\prime}\subset nch_{x}$, the CI test $X\stackrel{{\scriptstyle?}}{{\perp\!\!\!\perp}}Y|S^{\prime}$ is performed in C-PC. Thus the conditional independence $X\perp\!\!\!\perp Y|mns_{X}(Y)$ will be found and $X,Y$ are non-adjacent in $\hat{G}$, too. Second, we have to show that $\hat{G}$ and $\bar{G}$ have the same arrowheads. Due to the same skeleton, it is clear that they will also have the same unshielded colliders and same orientations due to Meek’s edge orientation rules. The only thing left to show that a) additional edge orientations coming from the C-DAG edges $E_{C}$ are the same in $\hat{G}$ and $\bar{G}$, as well as that b) edge orientations from using Meek’s orientation rules on the edges that partly come from a) are the same. Any directed edge $X\rightarrow Y\in E_{\bar{G}}$ coming from $B$ will also be directed in $\hat{G}$ due to Algorithm˜1. This then also leads to any directed edge $X\rightarrow Y\in E_{\bar{G}}$ coming from using orientation rules when combining CPDAG $G^{*}$ with $B$, also being oriented in the last steps of C-PC when Meek’s orientation rules are applied, so $X\rightarrow Y\in E_{\hat{G}}$. $\square$
A.4 Soundness of C-FCI, informativeness vs FCITiers
Theorem 5 (Soundness of Cluster-FCI).
If the C-DAG $G_{C}$ is compatible with ground truth MAG $G_{M}$, C-FCI is sound in the sense that an edge between any nodes $X_{i},X_{j}$ is present in the PAG $G_{P}$ output from C-FCI if and only if it is present in the ground truth MAG $G_{M}$. Any arrow or tail edge mark in $G_{P}$ is also present in $G_{M}$.
Proof: Cluster-FCI is the same as FCI, the only difference is it uses Algorithm˜3 as an ’oracle’ pre-processing step. As C-DAG $G_{C}$ is compatible with $G_{M}$, and any nodes $X_{i},X_{j}$ that are potentially connected by an inducing path are connected during Algorithm˜3, using this Algorithm does not remove any edges that FCI would not also remove. With the same reasoning, any arrowhead present in the partial ancestral graph $G_{P}$ is also present in $G_{M}$.
Remark 1 (Sketch for proof: C-FCI at least as informative as FCITiers).
Showing that C-FCI is at least as informative as FCITiers could be done as follows:
  1. (i)
    Construct tiers from C-ADMG $G_{C}$ by combining clusters, so that the tiers are TBK. (Otherwise, one can not compare; FCITiers can only work on TBK, not on arbitrary C-ADMGs)
  2. (ii)
    Show that any arrowhead oriented by FCITiers running on the previous TBK would also be oriented by running C-FCI on the C-DAG.
As the C-ADMG can contain bidirected edges, we can transform the C-ADMG to TBK as follows:
  1. (i)
    Order clusters $C_{1},\ldots,C_{r}$ topologically.
  2. (ii)
    Group all clusters that are connected by a bidirected path together into one tier $T_{i}$.
  3. (iii)
    The new cluster graph (now a C-DAG, without bidirected edges) could contain cycles. For any cycles in the C-DAG, merge the clusters on a cycle together into the same tier.
  4. (iv)
    Sort the tiers topologically.
  5. (v)
    Now one has a C-DAG that satisfies TBK.
C-FCI and FCITiers orient edge marks using the same collider rules and orientation rules. In addition, any extra edge marks in FCITiers come from two nodes $X,Y$ being in different tiers, e.g., $X\rightarrow Y$. But by construction, $X,Y$ were also in different clusters and if $X\rightarrow Y$ in TBK, then also $X\rightarrow Y$ in the partial mixed graph $G_{pm}$ from which C-FCI will start from. As again they use the same orientation rules, any arrow or tail edge mark from FCITiers will also be returned by C-FCI.
Appendix B Supplement to the simulation studies
B.1 Metrics for simulation studies
Definition 24 (Precision, recall, F1-score).
The precision is defined as
$precision=\frac{TP}{TP+FP},$ (2)
where TP = true positives and FP = false positives. Positive means the corresponding edge is present and negative means it is absent.
The recall is defined as
$recall=\frac{TP}{TP+FN},$ (3)
where FN = false negative, i.e., an edge was erroneously deleted. The F1-score is the harmonic mean of recall and precision and encourages balance between the two, as it is zero whenever one of them is zero,
$\text{F1-score}=\frac{precision*recall}{precision+recall}.$ (4)
Definition 25 (Structural Hamming distance).
The structural Hamming distance (SHD) between graphs $G,G^{\prime}$ is the number of edge deletions, additions or flips needed to transform $G$ into $G^{\prime}$.
Remark 2 (How arrow precision and recall are calculated).
Adjacency true/false positive/negative is easy to understand. For arrow marks, positive/ negative refers to arrow edge marks, so positive means arrow is there, negative means arrow is not there (tail/ circle edge mark). For example, a false positive is there if the true MAG says there is a circle edge mark, but the algorithm output says there is an arrow edge mark (at some edge between some nodes).
B.2 Parameters and additional tables from the simulation studies
See Tables˜3, 4 and 5.
Hyperparameter Simulation 1 Simulation 2 Simulation 3 Simulation 4
Algorithms PC, C-PC PC, C-PC FCI, FCITiers, FCI, FCITiers,
C-FCI C-FCI
Total number of graphs 1750 1080 180 180
Runs per configuration 10 1 10 5
DAG generation method Erdos-Renyi Erdos-Renyi, Erdos-Renyi Erdos-Renyi
hierarchical
scale-free
Distribution Gaussian Exponential, Gaussian Gaussian
Gaussian,
Gumbel
Alpha for CI test [0.01, 0.05, 0.1, [0.01, 0.05, 0.1, 0.05 0.05
0.25, 0.5] 0.25, 0.5]
CI test Fisher-z Fisher-z Fisher-z Fisher-z
Number of nodes 15 15 18 15
Number of edges [15, 30, 50, 80, 150] [15, 30, 50, 80] [18, 24, 30] [15, 20, 25]
Number of clusters [1, 2, 3, 4, 5, 6, 7] [1, 2, 3, 4, 5, 6] [2, 3, 4, 5, 6, 7] [3, 4, 5]
Sample size 1000 1000 [1000] [1000]
Weight range (-1, 2) (-1, 2) (-1, 2) (-1, 2)
Cluster method dag (Section˜B.3) dag dag cdag
Table 3: Hyperparameters for simulation studies 1–4
Table 4: Simulation 2: Base PC metrics for different distributions and DAG generation methods. The Base PC algorithm shows higher adjacency precision but generally lower recall and F1-score compared to Cluster-PC, a trend similar to Simulation 1. The structural Hamming distance (SHD) reflects similar trends, with performance depending on the underlying DAG structure.
Distribution Adj. precision Adj. recall Adj. F1-score SHD
Exponential 85.885.8% 46.546.5% 55.755.7% $40.3$
Gaussian 87.087.0% 46.546.5% 56.256.2% $40.1$
Gumbel 86.686.6% 46.246.2% 55.855.8% $40.4$
DAG method Adj. precision Adj. recall Adj. F1-score SHD
Erdos-Renyi 84.984.9% 54.554.5% 61.861.8% $32.1$
Hierarchical 90.390.3% 24.824.8% 38.738.7% $65.6$
Scale free 84.284.2% 60.060.0% 67.367.3% $23.1$
Table 5: Simulation 2: Cluster-PC metrics for different distributions and DAG generation methods. Cluster-PC typically shows slightly reduced adjacency precision compared to Base PC, but higher recall, better F1-scores, and improved structural Hamming distance (SHD), following the same patterns observed in Simulation 1.
Distribution Adj. precision Adj. recall Adj. F1-score SHD
Exponential 82.882.8% 56.456.4% 63.563.5% $34.3$
Gaussian 83.783.7% 56.456.4% 63.863.8% $33.9$
Gumbel 83.183.1% 55.755.7% 63.163.1% $34.6$
DAG method Adj. precision Adj. recall Adj. F1-score SHD
Erdos-Renyi 80.980.9% 63.563.5% 67.667.6% $27.4$
Hierarchical 89.389.3% 36.736.7% 51.651.6% $56.0$
Scale free 79.679.6% 68.468.4% 71.371.3% $19.5$
B.3 How compatible C-DAGs are generated
We called the method we used for the simulation studies 1-3 ’dag-first’. In this case we first generate a DAG and afterwards create a clustering by slicing up the topological ordering into $n$ cluster slices of random size.
For example, if the DAG has ten nodes and the number of clusters is three, this method will select two numbers between [1,n_clusters][1,n\_clusters], say four and ten. This means the first cluster will include the first three nodes in the topologiacl ordering, the second cluster contains nodes four to nine and the third cluster contains node ten.
For Simulation 4 we use another method we call ’cdag-first’. This is because we wanted to enforce TBK on the cluster graphs. This method first generates an Erdos-Renyi graph for the clusters, for example of size three again. Then it generates nodes for each cluster so that they sum up to the desired node number, say ten again. Then the graph is built according to the generated cluster graph and nodes, and some edges from that graph are dropped out, that probability is influenced by the n_edges parameter. Since the C-DAG is generated first, we can exclude bidirected edges. And a fully connected version of this C-DAG does satisfy TBK.
Appendix C Pairwise characterization of C-DAGs
A C-DAG can be represented as a boolean combination of pairwise background knowledge due to the following theorem:
Theorem 6 (Pairwise characterization of C-DAGs).
Clusters $C_{i},C_{j}$ imply, depending on their relationship in C-DAG $G_{C}$, pairwise constraints in the following ways:
  1. (i)
    $C_{i}=C_{j}\Rightarrow\textbf{T}$
    (the tautology, which is always true and places no restriction)
  2. (ii)
    $C_{i}\rightarrow C_{j}\Rightarrow\bigwedge_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\cancel{\dashleftarrow}X_{j}\wedge\bigvee_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\rightarrow X_{j}=:dir(C_{i},C_{j})$ (5)
    (note that $X_{i}\cancel{\dashleftarrow}X_{j}$ also implies $X_{i}\not\leftarrow X_{j}$)
  3. (iii)
    $C_{i}\dashrightarrow C_{j}\wedge C_{i}\not\rightarrow C_{j}\Rightarrow\bigwedge_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\cancel{\dashleftarrow}X_{j}\wedge X_{i}\not\rightarrow X_{j}=:anc(C_{i},C_{j})$ (6)
  4. (iv)
    $C_{i}\cancel{\dashrightarrow}C_{j}\wedge C_{i}\cancel{\dashleftarrow}C_{j}\wedge C_{i}\neq C_{j}\Rightarrow\bigwedge_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\cancel{\dashleftarrow}X_{j}\wedge X_{i}\cancel{\dashrightarrow}X_{j}=:nrel(C_{i},C_{j})$ (7)
These four points exhaust all possibilities in which two clusters could relate to each other in a C-DAG. The background knowledge implied by C-DAG $G_{C}$ with clustering $C=\{C_{1},\ldots,C_{m}\}$ can therefore be represented as a boolean combination in pairwise form:
$bk(G_{C}):=\bigwedge_{C_{i},C_{j}\in C}(\bigwedge_{C_{i}\rightarrow C_{j}}dir(C_{i},C_{j})\wedge\bigwedge_{\begin{subarray}{c}C_{i}\dashrightarrow C_{j}\\ C_{i}\not\rightarrow C_{j}\end{subarray}}anc(C_{i},C_{j})\wedge\bigwedge_{\begin{subarray}{c}C_{i}\neq C_{j},C_{i}\cancel{\dashrightarrow}C_{j}\\ C_{i}\cancel{\dashleftarrow}C_{j}\end{subarray}}nrel(C_{i},C_{j})).$ (8)
Thus the following equivalence holds:
$\text{$G$ compatible with $G_{C}$}\Longleftrightarrow\text{$bk(G_{C})$ is true for $G$}.$ (9)
Proof: (i): For two nodes in the same cluster $C_{i}$, no restriction is made, so for all $X_{i},X_{j}\in C_{i}$, T is true.
(ii): For two clusters $C_{i},C_{j}$ with $C_{i}\rightarrow C_{j}$ and nodes $X_{i}\in C_{i},X_{j}\in C_{j}$, it is impossible to have $X_{i}\dashleftarrow X_{j}$, as that would mean $C_{i}\dashleftarrow C_{j}$ in $G_{C}$, which is a cycle together with $C_{i}\rightarrow C_{j}$ in contradiction to $G_{C}$ being a C-DAG and $C$ being admissible. In addition, at least one $X_{i}\rightarrow X_{j}$ needs to be present by C-DAG construction. Therefore $C_{i}\rightarrow C_{j}$ implies $\bigwedge_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\cancel{\dashleftarrow}X_{j}\wedge\bigvee_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\rightarrow X_{j}$.
(iii): Let two clusters $C_{i},C_{j}$ have $C_{i}\dashrightarrow C_{j}\wedge C_{i}\not\rightarrow C_{j}$, i.e., there is a directed path from one to another, but they are not adjacent. With nodes $X_{i}\in C_{i},X_{j}\in C_{j}$ for analogous reasoning to (ii), it is impossible to have $X_{i}\dashleftarrow X_{j}$. In addition, as $C_{i}\not\rightarrow C_{j}$, it is also impossible to have $X_{i}\rightarrow X_{j}$. So $C_{i}\dashrightarrow C_{j}\wedge C_{i}\not\rightarrow C_{j}$ implies $X_{i}\cancel{\dashleftarrow}X_{j}\wedge X_{i}\not\leftarrow X_{j}$.
(iv): For two clusters that are not the same, not adjacent and not connected by a directed path, i.e., $C_{i}\cancel{\dashrightarrow}C_{j}\wedge C_{i}\cancel{\dashleftarrow}C_{j}\wedge C_{i}\neq C_{j}$ and nodes $X_{i}\in C_{i},X_{j}\in C_{j}$ it is impossible for $X_{i},X_{j}$ to be connected by a directed path. Therefore this implies $\bigwedge_{X_{i}\in C_{i},X_{j}\in C_{j}}X_{i}\cancel{\dashleftarrow}X_{j}\wedge X_{i}\cancel{\dashrightarrow}X_{j}$.
Equivalence statement: $\Rightarrow"$: Let $G=(V,E)$ be a graph over the same variables $V$ with edges $E$ being compatible with C-DAG $G_{C}$. Furthermore let $bk(G_{C})$ be as defined in Eq.˜8. The task is to show that $bk(G_{C})$ is true for $G$, specifically, that any edge between any $X_{i},X_{j}$ satisfies $bk(G_{C})$. Take any $X_{i},X_{j}\in V$ and their respective clusters $X_{i}\in C_{i},X_{j}\in C_{j}$. $C_{i},C_{j}$ relate to each other in exactly one of the four ways described in (i)-(iv). Whatever the cluster relationship is on the left hand side of (i)-(iv), the proofs of (i)-(iv) above show that the edge between $X_{i},X_{j}$ satisfies the constraint on the right hand side of (i)-(iv). So the boolean pairwise combination $bk(G_{C})$ is true for $G$.
$\Leftarrow"$: Let $G_{C}$ be a C-DAG and $bk(G_{C})$ be true for DAG $G$. The task is to show that $G$ is compatible with $G_{C}$. $G=(V,E)$ is compatible with $G_{C}$, if none of its edges $E$ contradict the C-DAG edges $E_{C}$. Take any $X_{i},X_{j}$ and their corresponding clusters $X_{i}\in C_{i},X_{j}\in C_{j}$. The edge between $X_{i},X_{j}$ can take any of the three forms $X_{i}\rightarrow X_{j}$, $X_{i}\leftarrow X_{j}$ or $X_{i}\cancel{-}X_{j}$. W.l.o.g. (no need to consider the case $C_{i}\leftarrow C_{j}$ due to symmetry) the C-DAG restriction on the edge between $X_{i},X_{j}$ can take any of the forms in (i)-(iv).
If restriction T is put on the edge between $X_{i},X_{j}$, they are put in the same cluster and any of $X_{i}\rightarrow X_{j}$, $X_{i}\leftarrow X_{j}$ and $X_{i}\cancel{-}X_{j}$ would be compatible with $G_{C}$. If restriction $X_{i}\cancel{\dashleftarrow}X_{j}$ is put, $X_{i}$ and $X_{j}$ are in adjacent clusters $C_{i}\rightarrow C_{j}$. The edges allowed by $X_{i}\cancel{\dashleftarrow}X_{j}$ are $X_{i}\rightarrow X_{j}$ and $X_{i}\cancel{-}X_{j}$. Both of them are compatible with $G_{C}$. If restriction $X_{i}\cancel{\dashleftarrow}X_{j}\wedge X_{i}\not\rightarrow X_{j}$ or $X_{i}\cancel{\dashleftarrow}X_{j}\wedge X_{i}\cancel{\dashrightarrow}X_{j}$ is put, only $X_{i}\cancel{-}X_{j}$ is allowed and $C_{i},C_{j}$ are not adjacent, so $X_{i}\cancel{-}X_{j}$ is compatible with $G_{C}$. $\square$
Theorem 7 (Pairwise characterization of C-ADMGs).
Let the setup be the same as in Theorem˜6. To include bidirected constraints, the rules (i)-(iv) from Theorem˜6 get extended by
  1. (v)
    $C_{i}\not\leftrightarrow C_{j}\Longleftrightarrow\bigwedge_{\begin{subarray}{c}X_{i}\in C_{i},X_{j}\in C_{j}\end{subarray}}X_{i}\not\leftrightarrow X_{j}=:nlat(C_{i},C_{j})$
The background knowledge is then denoted as
$bk(G_{C})=\bigwedge_{C_{i},C_{j}}(\bigwedge_{C_{i}\rightarrow C_{j}}dir(C_{i},C_{j})\wedge\bigwedge_{\begin{subarray}{c}C_{i}\dashrightarrow C_{j}\\ C_{i}\not\rightarrow C_{j}\end{subarray}}anc(C_{i},C_{j})\\ \wedge\bigwedge_{\begin{subarray}{c}C_{i}\cancel{\dashrightarrow}C_{j}\\ C_{i}\cancel{\dashleftarrow}C_{j}\end{subarray}}nrel(C_{i},C_{j})\wedge\bigwedge_{\begin{subarray}{c}C_{i}\not\leftrightarrow C_{j}\end{subarray}}nlat(C_{i},C_{j}))$ (10)
and
$\text{$G$ compatible with $G_{C}$}\Longleftrightarrow\text{$bk(G_{C})$ is true in $G$}$ (11)
Proof: (v) If $C_{i}\not\leftrightarrow C_{j}$, by C-ADMG definition that means for all $X_{i}\in C_{i},X_{j}\in C_{j}$ it is $X_{i}\not\leftrightarrow X_{j}$ which exactly implies $nlat(C_{i},C_{j})$. For the reverse direction, if $nlat(C_{i},C_{j})$ it means for no $X_{i}\in C_{i},X_{j}\in C_{j}$ it is $X_{i}\leftrightarrow X_{j}$. This means the C-ADMG does not have $C_{i}\leftrightarrow C_{j}$. The rest follows analogously to the proof of Theorem˜6. $\square$
Appendix D Using C-DAGs for score-based and continuous optimization discovery algorithms
In this paper we studied constraint based causal discovery, but DAGs can also be estimated via a constrained optimization problem (chickering2002optimal; zheng2018dags). Such a problem typically admits a form like
$\min_{G\in\mathbb{R}^{d\times d}}F(G)\quad\text{subject to}\quad G\in\text{DAGs}$
with $F(G)$ evaluating the goodness of fit of a graph $G$ to the available data. One can easily extend this problem to include C-DAG constraints:
$\min_{G\in\mathbb{R}^{d\times d}}F(G)\quad\text{subject to}\quad G\in\text{DAGs},G\text{ compatible with }G_{C}$
so that the optimization procedure is only able to take steps in the space of DAGs that are compatible with C-DAG $G_{C}$ or only able to return optimal solutions that are compatible with $G_{C}$. This could be an interesting topic for future research.